(4π‘₯𝑦 + 3𝑦2 )𝑑π‘₯ + π‘₯(π‘₯ + 2𝑦)𝑑𝑦 = 0


1
Expert's answer
2021-10-12T13:35:39-0400

"(4xy+3y^2)_y=4x+6y\neq (x(x+2y))_x=2x+2y"

So, the equation is not exact.


"u=y\/x"

"y'=xu'+u"

"3x^2u^2+2x^2(xu)'+4x^2u+x^2(xu)'=0"

"2x^3uu'+x^3u'+5x^2u^2+5x^2u=0"


"\frac{(2u+1)u'}{5u(u+1)}=-\frac{1}{x}"


"\int\frac{2u+1}{5u(u+1)}du=-\int\frac{1}{x}dx"


"\frac{ln(u^2+u)}{5}=c-lnx"


"u_1=-\frac{\sqrt{c_1/x+1}+1}{2}"


"u_2=\frac{\sqrt{c_1/x+1}-1}{2}"


"y_1=-\frac{\sqrt{c_1/x+1}+1}{2}x"


"y_2=\frac{\sqrt{c_1/x+1}-1}{2}x"